OpenStudy (anonymous):

A bird flies from the bottom of the cave during the time period 0

OpenStudy (anonymous):

Is the equation: \[\sqrt{(t)-1}\]?

OpenStudy (anonymous):

\[v(t) = \frac{dx}{dt}\]

OpenStudy (anonymous):

\[dx = v(t) dt\]

OpenStudy (anonymous):

\[\int_0^x dx = \int_0^4v(t) dt\]

OpenStudy (anonymous):

no no sorry just the square root of t

OpenStudy (anonymous):

then minus one

OpenStudy (anonymous):

Now I am going to solve the RHS(Right Hand Side)

OpenStudy (anonymous):

\[v(t) = \sqrt t -1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[RHS = \int_0^4 \sqrt t dt- dt\]

OpenStudy (anonymous):

\[RHS = \frac{2}{3} t^{3/2} - t |_0^4 \]

OpenStudy (anonymous):

Then On whole it becomes \[x = \frac{16}{3} - 4\]

OpenStudy (anonymous):

Its positive?

OpenStudy (amistre64):

distance traveled would correlate to a line integral \[\int\sqrt{[f'(x)]^2+[g'(x)]^2}\]

OpenStudy (anonymous):

i know it sounds stupid but my first time i got a negative number. I guess it was just another silly math mistake :)

OpenStudy (anonymous):

Ok I got Displacement Not Distance

OpenStudy (anonymous):

I don't know about the distance but displacement should be the one I got

OpenStudy (anonymous):

ohhh no.

OpenStudy (anonymous):

And Distance is always positive

OpenStudy (amistre64):

since there is only 1 equation to deal with here; i believe it goes: \[\int_{a}^{b}\ \sqrt{1+\left(\frac{1}{2\sqrt{t-1}}\right)^2}\ dt\]

OpenStudy (anonymous):

You can't have negative

OpenStudy (anonymous):

So if we have the displacement, how can we use that to find distance?

OpenStudy (anonymous):

Distance Traveled= \[\int\limits_{0}^{4} (\sqrt{(t)}-1) dx\] =(t)^(1/2)-1 =(t)^(3/2)/(3/2) - (t) =(2/3)(t^(3/2)) - (t) Now plug in t=4 solve then t=0 solve. Subtract the result of t=4 from result of t=0 Distance=4/3

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

Thats just what I got justtheretohelp but it is Displacement

hero (hero):

The integral of velocity is distance, right?

hero (hero):

Oh, wait....

hero (hero):

Yeah, it's displacement

hero (hero):

lol, I agree

OpenStudy (anonymous):

Well the integral of the absolute value of the v(t) function from b to a would be distance

OpenStudy (amistre64):

but then i believe you stated its sqrt(t) - 1 \[\int_{a}^{b}\ \sqrt{1+\left(\frac{1}{2\sqrt{t}}\right)^2}\ dt\] \[\int_{a}^{b}\ \sqrt{1+\left(\frac{1}{4t}\right)}\ dt\] \[\int_{a}^{b}\ \sqrt{\frac{4t+1}{4t}}\ dt\] \[\int_{a}^{b}\ {\frac{\sqrt{4t+1}}{2\sqrt{t}}}\ dt\] have a messed it up yet :)

OpenStudy (anonymous):

not to put in my two cents at the last minute, but what do you make of this line . If the bird moves along a curve, so that its velocity at time t is v(t)= the square root of (t) -1

OpenStudy (anonymous):

you are given the velocity, but not the curve!

OpenStudy (amistre64):

...great, now we need to rescript the whole thing. cut!!

hero (hero):

bravo!

OpenStudy (anonymous):

you have \[v(t)=\sqrt{t}-1\] and so if the bird was moving in a straight line, then you would take the integral over positive and negative part of this and add, but we have no idea what the bird is doing do we?

OpenStudy (anonymous):

of course i could be wrong.

OpenStudy (amistre64):

if theres an option to click "next"... take it now :)

OpenStudy (anonymous):

@amistre you are finding the length of the curve \[y=\sqrt{t}-1\] from 0 to 4

OpenStudy (amistre64):

that i was .... that i was

hero (hero):

lol

OpenStudy (anonymous):

now from 0 to 1 apparently the bird is flying backwards

OpenStudy (anonymous):

maybe it is a hummingbird in a cave.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Gee.. I always thought distance and displacement were the same thing.. :/

OpenStudy (anonymous):

i will make a guess. my guess is your teacher wants \[\int_0^1 1-\sqrt{t}dt +\int_1^4\sqrt{t}-1 dt\]

OpenStudy (anonymous):

no displacement is not the same as distance. if you go back 4 units and forward 6 your displacement is 2 and your distance is 10

OpenStudy (anonymous):

that is you traveled ten miles but are only two miles from home

OpenStudy (anonymous):

So how do we get to the distance

OpenStudy (anonymous):

integrate over where it is positive and negative separately. then add

OpenStudy (anonymous):

Agree, with satellite73. Didn't think it through the first time.

OpenStudy (anonymous):

i should say integrate over where it is negative, then take absolute value, then add

OpenStudy (amistre64):

could we integrate up to a position function?

OpenStudy (anonymous):

Gosh, and I thought mere integration was complex...

OpenStudy (amistre64):

then the derivative of the position function would amount to the velocity function to use in the distance formula?

OpenStudy (amistre64):

in other words, we already have h'(x) given right?

OpenStudy (anonymous):

\[\int_0^1\sqrt{t}-1dt=-\frac{1}{3},\int_1^4 \sqrt{t}-1 dt =\frac{5}{3}\] so i will bet the answer is \[\frac{1}{3}+\frac{5}{3}=2\]

OpenStudy (anonymous):

that is my bet, but i am not placing odds

OpenStudy (amistre64):

\[\int \sqrt{1+(\sqrt{t}-1)^2}\ dt\] maybe?

OpenStudy (anonymous):

@amistre i think what this means is we have velocity, just want distance traveled.

OpenStudy (anonymous):

\[v(t)=S'(t)=\sqrt{t}-1\] so \[S(t)=\frac{2}{3}t^{\frac{3}{2}}-t\]

OpenStudy (amistre64):

right, and in my twisted brain im thinking that the v(t) given is the derivative of the position function to place inside of the line integral as [V'(t)]^2?

OpenStudy (anonymous):

you are thinking too hard. it was that "curve" that threw you off. threw you a curve me might say

OpenStudy (amistre64):

:) yousa mighta be right

OpenStudy (amistre64):

if my thought is right, in some alternate universe as that may be :) ; I get this: http://www.wolframalpha.com/input/?i=integrate%28sqrt%28x-2sqrt%28x%29%2B2%29%29dx+from+0+to+4 abt 4.59 meters